OpenStudy (anonymous):

how to solve this guys. your help would be gladly appreciated :D log(base 2) 5+ log(base3) 5

OpenStudy (anonymous):

\[\log_2{5} + \log_3{5} ?\]

OpenStudy (tkhunny):

There is nothing to solve. That requires an equation of some sort.

OpenStudy (anonymous):

so this equation cant be solved?

OpenStudy (imstuck):

It can be simplified with the change of base formula. You can do this:\[\log _{2}(5)=\frac{ \log(5) }{\log(2) }\]and \[\log _{3}(5)=\frac{ \log(5) }{ \log(3) }\]

OpenStudy (imstuck):

\[\frac{ \log(5) }{\log(2) }=2.322\]and \[\frac{ \log(5) }{\log(3) }=1.465\]If you add those together you get 3.787. That's all I can come up with for this.

OpenStudy (anonymous):

okay. I got it. thanks for the help :D

OpenStudy (imstuck):

Are you using the change of base formula in math class?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

I just wanted tp be sure if my answer was correct.

OpenStudy (tkhunny):

It isn't an equation. And thus, it cannot be solved. The expression can be evaluated. Next time you want to check your answer, first SHOW your work.